Suppose that the MPC is .80 and government spending increases by 500. Using the simple multiplier, output would change by:

A) 2000

B) 2500

C) 400

D) 1600
Answer #1

B is right option that output would change by 2500

Simple multiple times formula= 1/(1- MPC)

= 1/(1-0.80)

=5

So$ 500 intial increase in investment= 5×500

= 2500